# Automatically Analyzing Performance Issues in Android Apps: How Far Are We?

### v1

@@ 1. Introduction
However, due to the non-functional nature of performance issues, their problems do not immediately manifest. Additionally, performance issues can be caused by multiple factors, requiring in-depth analysis and debugging to diagnose and identify the root causes (Liu2016FixingRL, ; Habchi2018OnAL, ; Jabbarvand2019SearchBasedET, ; Li2019CharacterizingAD, ; Liu2014GreenDroidAD, ). Furthermore, some performance issues only appear under specific usage scenarios or loads (Behrouz2016EnergyawareTM, ; Jabbarvand2019SearchBasedET, ; Yan2013SystematicTF, ; Farooq2018RuntimeDroidRR, ). Consequently, manually identifying performance issues is time-consuming and labor-intensive, making it difficult for developers to address all problems during the testing phase. Therefore, it is crucial to use effective automated tools to analyze, detect, and fix performance issues (Habchi2018OnAL, ; Li2019DetectingAD, ; Li2022CombattingEI, ).
Unfortunately, at the current stage, researchers face difficulties in designing automated tools to comprehensively address Android performance issues due to: 1) Lack of clear classification of performance issues and their root causes, making it hard to fully understand the performance issues; 2) Uncertainty about the state quo of existing approaches, making it unclear to what extent these tools are reliable for detecting or fixing performance issues; 3) The absence of benchmark datasets in evaluating state-of-the-art tools; 4) Uncertainty about whether state-of-the-art tools and datasets accurately reflect real performance issues.
To guide future research, we first conducted a systematic literature review to provide a holistic taxonomy for existing performance issues and their underlying root causes. We then identified research gaps between state-of-the-art tools and datasets relative to the targeted research objectives (i.e., performance issues). Subsequently, we analyzed real-world performance issues to pinpoint differences between existing literature and practical challenges. Through these investigations, our goal is to offer a thorough understanding of Android performance issues and to guide the development of more effective automated tools and comprehensive datasets for evaluation. In this paper, we aim to explore the following research questions:
RQ1: What is the status quo of Android performance issues analysis in research field?
  • What performance issues are being targeted by researchers? The lack of a universally accepted taxonomy leads to many papers targeting the same performance problems using different terms. Additionally, since one factor can lead to multiple performance issues and one issue can be caused by multiple factors, creating a clear classification system is challenging. Our work filled this gap by distinguishing between the consequences and factors of performance issues. We identified five categories of Android performance issues and 21 distinct root causes.
  • What approaches have been developed for performance issues? To solve the above performance issues, researchers have proposed various kinds of techniques. Thus, in the second research question, we aim to understand the status quo of current techniques, seeking to comprehensively grasp the research gap between performance issues and automatic tools. In total, we classified them into three primary techniques: static-based, dynamic-based, and deep learning-based approaches. Static analysis is the most frequently adopted technique, while deep learning-based methods are much less commonly utilized for performance issues. Moreover, it is surprising to find that only 7 out of 21 factors are targeted by these tools. This evidence confirms a significant gap between performance issues and automated resolutions. Additionally, the majority of these tools focus on issue detection rather than fixing, highlighting another research gap concerning the need for proactive problem-solving.
  • What datasets have been used in Android performance analysis? High-quality datasets are essential for evaluating performance analysis tools. We found that current datasets cover only 11 out of 21 factors, addressing a limited range of issues. Furthermore, the number of existing datasets is very small. This highlights an urgent need for a comprehensive public dataset.
RQ2: What are gaps between cutting-edge studies and real-world scenarios?
  • What real-world performance issues in the wild? Given that performance is an inherently practical concern, we proceeded to conduct a deeper analysis of real-world issues to pinpoint differences between existing literature and practical challenges. Our examination revealed one previously unrecognized performance issue and 19 previously unacknowledged factors across real-world platforms. This suggests that numerous factors contributing to performance issues in practical settings have not received attention in research. Furthermore, we uncovered a misalignment between the primary concerns of researchers, developers, and users. These discoveries underscore a substantial gap in fully comprehending and mitigating performance issues in real-world scenarios.
  • What are the common patterns exhibited in real-world performance issues? Based on our systematic analysis for real-world data, we identified six common code patterns: API Misuse, Unreleased References, Redundant Objects, Large-Scale Data, UI Operations, and Other Patterns. These newly discovered patterns provide insights for designing automated tools to address performance issues and for creating datasets to evaluate the effectiveness of these tools.
2. Studies on Performance Issues
In this work, we are interested in how far we are towards automating performance issues analysis in Android apps. Thus, in this section, we first resort to a systematic literature review (SLR) to understand the status quo of Android app performance analysis.
2.1. Literature Review
To conduct systematic literature review, we followed the methodologies outlined by Brereton et al.  (brereton2007lessons, ) and Kitchenham et al.  (Kitchenham2004ProceduresFP, ). Fig. 1 illustrates the working process which comprises 5 steps.
2.1.1. Keywords Identification
To investigate the status quo of Android performance analysis, we identified a bunch of search keywords to locate relevant publications on this topic. Specifically, the keywords comprise two groups (i.e., G1 AND G2). Each group is made up of several keywords. The search string is then constructed by combining them, i.e., g1 AND g2, where g1 and g2 represent the disjunction of keywords from groups G1 and G2, respectively. Please note that these keywords are based on our initial understanding of performance issues and are intended to help us identify “seed” papers. As we review these papers, we could uncover new performance issues, prompting a snowballing process that could lead us to further relevant studies. The * symbol represents any string of characters. For example, *phone* can refer to smartphone, phones, or phone. Similarly, issue* can represent issues or issue. Keywords are as follows:
G1: android, mobile, *phone*
G2: performance, resource, energy, responsiveness, issue*
Figure 1. The working process of our systematic literature review.
Table 1. CORE A/A* ranked software engineering and mobile venues.
Type Source Research Codes Venues
Journals CORE2020 0803, 0805 TSE, TOSEM, JSS, IST, TMC
Conferences CORE2023 4606, 4612 MOBICOM, Mobisys, EuroSys, MOBIHOC, MSWIM, ASE, ESEC/FSE, ICSE, ISCA, EASE, ECSA, ESEM, FOSSACS, ICSA, ICSME, ICST, ISSRE, ISSTA, MSR, SANER, SEAMS
2.1.2. Repository Search
To ensure systematic literature collection, we conducted keyword-based searches across five libraries: IEEE Library, ACM Library, Springer Library, Science Direct Library, and Wiley Library.
From this initial search, we retrieved a total of 10,766 papers. However, not all of them were of high quality, and many were outside our area of focus. To narrow down to the most relevant papers, we concentrated on the fields of software engineering and mobile, selecting research codes 4606 and 4612 for conferences, and 0805 and 0803 for journals, specifically targeting publications containing the keywords “software” and “mobile”. In addition, to further guarantee the quality of papers, we exclusively chose those published from conferences rated CORE A*/A  [^1] CORE provides conference and journal ratings. https://www.core.edu.au/home. We then systematically reviewed each of these venues, applying the aforementioned keywords to search for the title of relevant publications. As a result, we successfully identified 274 publications distributed across 22 venues as listed in Table 1 (i.e., there is no relevant paper identified in 4 of the venues).
2.1.3. Paper Exclusion
To gather relevant papers as possible, we initially considered all returned results from our search. However, not every paper pertains to Android performance issues. To refine our collection, we conducted a thorough review of the abstracts (and full content when necessary) of these papers, aiming to retain only those closely related to our research focus. Specifically, we applied the following exclusion criteria to streamline this process:
  1. (1)
  2. (2)
    Papers targeting on non-Android mobile devices (e.g., Windows Phone, IOS devices) were excluded.
  3. (3)
    Papers using Android performance issues as metrics/indicators for comparison were excluded. For example, Mantis (Kwon2013MantisAP, ) predicts app execution time based on app implementations.
  4. (4)
    Papers targeting on Android performance issues but not focused on the mobile app level were excluded. For example, Hypnos (Jindal2013HypnosUA, ) is used for detecting sleep conflicts in smartphone device drivers, which is outside our scope focused on mobile app issues stemming from poor coding practices by developers.
After applying these exclusion criteria, we retained 27 papers that are closely related to the topic of performance issues analysis at Android mobile level.
2.1.4. Back/Forward Snowballing
After initially identifying relevant papers (i.e., 27 papers), we employed a backward and forward snowballing method to ensure comprehensive coverage of closely related literature. The backward snowballing process involved scrutinizing the references of identified papers for additional relevant sources. Additionally, we performed the forward snowballing method by examining papers citing the identified papers. We strictly adhered to the exclusion criteria when conducting snowballing until no further papers could be identified. Eventually, we were able to locate a total of 48 papers. Please note that during snowballing, we also collected 82 keywords related to performance issues for real-world data collections in Section 3.1.2.
2.1.5. Cross Check
Even with the relevant selection criteria, it is not guaranteed that the quality of all identified papers is desirable. Therefore, we evaluated the quality of each paper by reading its full text. Additionally, to avoid overlooking any primary studies, we conducted cross-checks involving at least two authors, ensuring the reliability of these papers and enhancing the trustworthiness of the primary studies. Ultimately, we collected 48 papers to answer our research questions defined in Section 1.
2.2. Result
In the following paragraphs of Section 2, we summarize the state quo from thorough reading of these papers and attempt to understand: 1) Targeted performance issues, which involve key factors that lead to problems needing resolution, 2) Automatic techniques designed to address mobile performance issues, and 3) Datasets designed to evaluate the effectiveness of these automatic techniques.
Table 2. The top problems targeted by the examined papers.
Category Factor Problem Example
Energy Consumption Database Operation Frequently triggering database queries for auto-complete (Li2019DetectingAD, ; Lyu2018RemoveRF, ; Li2022CombattingEI, ; Lyu2017AnES, )
Loop Executing energy-intensive behaviors in loop  (Banerjee2014DetectingEB, ; Behrouz2016EnergyawareTM, ; Lyu2018RemoveRF, ; Fratantonio2015CLAPPCL, ; Lyu2017AnES, )
HTTP Request Dividing a large packet into small pieces and transmitting data frequently (Li2016AutomatedEO, )
Image Decoding identical image in multiple times (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ; Carette2017InvestigatingTE, ; Kang2016ExperienceRD, )
Wakelock Forgetting to release wakelocks after usage  (Liu2014GreenDroidAD, ; Pathak2012WhatIK, ; Liu2016UnderstandingAD, ; Li2022CombattingEI, ; Couto2020EnergyRF, ; Li2020EnergyDxDE, ; Behrouz2016EnergyawareTM, ; Banerjee2018EnergyPatchRR, ; Banerjee2014DetectingEB, ; Pathak2012WhereIT, )
Resource Updating unnecessary sensor data in the background  (Liu2014GreenDroidAD, ; Liu2013WhereHM, ; Li2019DetectingAD, ; Behrouz2016EnergyawareTM, ; Li2022CombattingEI, ; Liu2016FixingRL, ; Couto2020EnergyRF, ; Jabbarvand2020AutomatedCO, ; Li2020EnergyDxDE, ; Guo2013CharacterizingAD, ; Lu2022DetectingRU, ; Banerjee2018EnergyPatchRR, ; Wu2016LightWeightIA, ; Jabbarvand2019SearchBasedET, ; Banerjee2014DetectingEB, ; Yan2013SystematicTF, )
UI Rendering Views are rendered but not visible to users because they are off-screen or set to invisible  (Gao2017EveryPC, ; Lin2024AgingOG, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, )
Inefficient Method Calling unnecessary methods (Carette2017InvestigatingTE, ; Couto2020EnergyRF, ; Habchi2019TheRO, )
View Drawing all items in a ListView without reusing exising data  (Habchi2019TheRO, ; Couto2020EnergyRF, )
HashMap Managing small collections with HashMap (Should use ArrayMap and SparseArray instead) (Couto2020EnergyRF, ; Habchi2019TheRO, ; Carette2017InvestigatingTE, )
Responsiveness Obsolete Task Downloading obsolete content occurs when switching activities  (Kang2016DiagDroidAP, ; Couto2020EnergyRF, )
UI Rendering Improperly stacking and layering in layouts can result in pixels being redrawn multiple times in the same frame, with only the last rendering visible (Gao2017EveryPC, ; Lin2024AgingOG, ; Liu2014CharacterizingAD, ; Habchi2019TheRO, )
Message Backloging rapidly incoming messages in Message Handler’s single-threaded (Kang2016DiagDroidAP, )
Main Thread Performing Heavy operations in the main thread  (Liu2014CharacterizingAD, ; Brocanelli2018HangDR, ; Lin2014RetrofittingCF, )
Activity Loading new resources during activity restarts may freeze the UI  (Farooq2018RuntimeDroidRR, ; Riganelli2020DataLD, ; Zhou2023DDLDroidED, ; Guo2022DetectingAF, ; Chen2023TransparentRC, ; Lu2022DetectingRU, )
View Allocating new objects along with draw operations (Habchi2019TheRO, ; Couto2020EnergyRF, )
Memory Consumption Obsolete Task Retaining obsolete parameters after updating view layouts  (Kang2016DiagDroidAP, ; Couto2020EnergyRF, )
Service Failing to destroy a dead service  (Song2019ServDroidDS, )
Resource Improper releasing of resources (e.g., Sensor, Camera)  (Liu2014GreenDroidAD, ; Liu2013WhereHM, ; Li2019DetectingAD, ; Behrouz2016EnergyawareTM, ; Li2022CombattingEI, ; Liu2016FixingRL, ; Couto2020EnergyRF, ; Jabbarvand2020AutomatedCO, ; Li2020EnergyDxDE, ; Lu2022DetectingRU, ; Banerjee2018EnergyPatchRR, ; Wu2016LightWeightIA, ; Jabbarvand2019SearchBasedET, ; Banerjee2014DetectingEB, ; Yan2013SystematicTF, )
HashMap Managing small collections with HashMap (Should use ArrayMap and SparseArray instead) (Couto2020EnergyRF, ; Habchi2019TheRO, ; Carette2017InvestigatingTE, )
Data Loss Activity Failing to save and restore GUI data during activity restarts from runtime changes (Farooq2018RuntimeDroidRR, ; Riganelli2020DataLD, ; Zhou2023DDLDroidED, ; Guo2022DetectingAF, ; Chen2023TransparentRC, ; Lu2022DetectingRU, )
AsyncTask Updating destroyed GUI components when it has been recreated using AsyncTask (Lin2015StudyAR, ; Pan2020StaticAC, ; Kang2016DiagDroidAP, ; Lin2014RetrofittingCF, ; Fan2018EfficientlyMA, )
Distorted UI Display Activity Overlapping fragments occur when a new activity is restarted with a new fragment attached while the old fragment is still retained by the system (Farooq2018RuntimeDroidRR, ; Riganelli2020DataLD, ; Zhou2023DDLDroidED, ; Guo2022DetectingAF, ; Chen2023TransparentRC, ; Lu2022DetectingRU, )
2.2.1. Result - Targeted Problem
We now highlight the representative problems targeted by our fellow researchers (cf. Table 2). Based on the literature review results, we manually summarized and identified five major categories of Android mobile performance issues: Energy Consumption, Memory Consumption, Responsiveness, Data Loss, and Distorted UI Display. Then we reviewed the root causes (a.k.a factors) for each category thoroughly and demonstrate them in Table 2 along with the specific performance problems our researcher fellows are working on.
Here, we remind readers that during the process of manual summarization, we found that the descriptions of performance issues are chaotic. The lack of a universally accepted taxonomy results in many papers targeting the same performance problems but using different descriptive terms. Additionally, since one factor can lead to multiple performance issues (consequences), and one consequence can be caused by multiple factors, it complicates the creation of a clear and comprehensible performance classification system. Therefore, our work addresses this gap by distinguishing between the consequences and factors of performance issues. Specifically, we first classify performance issues by their consequences, then define a series of factors that cause these issues, and finally link the consequences to their corresponding factors. To the best of our knowledge, we are the first to systematically classify all Android performance issues in a hierarchical manner. We now elaborate on the five major categories, each with the corresponding factors along with an example presented in Table 2.
Energy Consumption. As shown in Fig. 2, energy consumption receives the most research attention, with 34 out of the examined papers (70.83%) and its increasing trend is expected to continue. This result is not surprising given that the most urgent task at the moment is to ensure that the app remains user-friendly and sustainable in the rapidly evolving app market, which is why many previous studies focused on this research topic. Specifically, we found that most of the energy consumption issues stem from a common pattern: frequent energy-consuming operations occurring together with loops/lifecycle functions. For example, Lyu et al. (Lyu2018RemoveRF, ) found that database writes within loops trigger inefficient autocommit behaviors, leading to significant energy waste. Liu et al. (Liu2014CharacterizingAD, ) discovered that even when the GUI is in the background and not visible, threads continue updating the UI, causing unnecessary energy waste. In addition, invoking APIs for requiring resources (e.g., wakelock, sensor) when they are not needed can also results in a substantial increase in energy consumption (Liu2014GreenDroidAD, ; Li2022CombattingEI, ; Couto2020EnergyRF, ; Li2020EnergyDxDE, ; Banerjee2018EnergyPatchRR, ).
Responsiveness. Another important aspect of Android performance issues is responsiveness. Fig. 2 shows that 18 reviewed articles (37.50%) are crafted for poor responsiveness, leading to perceptible delays such as image rendering delays, UI freezes, unresponsiveness, etc. Specifically, we identified 10 distinct factors that contribute to this category. Most primary studies (Liu2014CharacterizingAD, ; Brocanelli2018HangDR, ; Lin2014RetrofittingCF, ) discovered that heavy operations associated with UI threads can slow responsiveness. Additionally, Gao et al. (Gao2017EveryPC, ) and Sarra et al. (Habchi2019TheRO, ) show that improper stacking and layering in layouts can cause pixels to be redrawn multiple times within the same frame, slowing down the response time.
Memory Consumption. There are 14 papers (29.17%) related to memory consumption issues in the collected studies. These memory consumption issues have diverse factors, often stemming from inefficient coding that leads to excessive memory use. For example, decoding images without resizing them results in extensive memory consumption (Song2021IMGDroidDI, ; Li2019CharacterizingAD, ). Additionally, failing to destroy an AsyncTask when it holds references to active GUI components can prevent the garbage collection of the destroyed GUI, leading to memory leaks (Lin2015StudyAR, ; Pan2020StaticAC, ).
Data Loss With regard to the data loss problem, Fig. 2 indicates that 7 papers (14.58%) attempt to analyze data loss issues, which significantly degrade user experience when data fails to be saved or restored in UI threads. For instance, studies have revealed that failing to save and restore GUI data during activity restarts can cause existing data to vanish instantly when the screen orientation changes (Farooq2018RuntimeDroidRR, ; Riganelli2020DataLD, ; Zhou2023DDLDroidED, ; Guo2022DetectingAF, ; Chen2023TransparentRC, ).
Distorted UI Display. Another aspect of performance issues relates to distorted UI display, which consists of only 2.08% (i.e., 1 study) of the collected studies. Specifically, Umar et al. (Farooq2018RuntimeDroidRR, ) shows that overlapping fragments occur when initializing a new fragment while retaining the old one, causing a poor user experience.
Figure 2. Annual publications on each performance issue.
Discussion. Among all the collected studies on performance issues for Android, it appears that energy consumption is the most frequently studied topic, followed by responsiveness analysis. Additionally, it is noteworthy that an increasing number of recent studies have started to address specific issues related to memory consumption and data loss (e.g., runtime change handling (Farooq2018RuntimeDroidRR, ; Chen2023TransparentRC, )), although the total number of relevant studies remains small. Moreover, the amount of literature focusing on distorted UI display is relatively small, indicating a need for further in-depth research. While this review categorizes the existing literature’s research objectives into five categories, the scope of Android performance issues can be much broader. Therefore, future research should go beyond these categories and propose new approaches tailored to emerging app scenarios.
Answers to RQ1.1: Current research studies focus on five categories of performance issues, prioritized by their prevalence: Energy Consumption (70.83%), Responsiveness (37.50%), Memory Consumption (29.17%), Data Loss (14.58%), and Distorted UI Display (2.08%). We provide a holistic taxonomy of Android performance issues, encompassing 5 major categories and 21 distinct root causes. Performance issues are sophisiticated, as a single issue may stem from multiple factors and a single factor can cause multiple issues.
2.2.2. Result - Technique
To address the above performance issues, researchers have proposed automatic tools using diverse techniques, categorized into static approaches, dynamic approaches, and deep learning approaches. Table 3 presents the the list of available automatic approaches collected from our examined papers. Despite numerous claims of proposed automated tools, only a small fraction (i.e., 10 out of the 48 examined papers (20.83%)) works have made their tool publicly available. This indicates a significant gap between automatic tools with their practical usage to within our community. In addition, among the identified factors leading to performance issues, only 7 out of 21 are targeted by existing tools. The fact that the majority of the factors are overlooked by SOTA tools indicates a huge gap in addressing performance issues. Moreover, most automatic tools are designed with a primary focus on simply detecting performance issues rather than actively fixing them. This emphasizes a need for a shift towards developing tools that not only identify but also offer solutions to enhance performance. To better understand how these available tools are implemented, we now summarize the techniques adopted in these approaches.
Tool/Reference Approach Technique Stage Target Factor Link
DDLDroid (Zhou2023DDLDroidED, ) Static Reverse Engineering Taint analysis Detect Activity  (DDLDroidTool, )
DLD (Riganelli2020DataLD, ) Dynamic Runtime Monitoring Model-based Testing Detect Activity  (Riganelli2020DataLDTool, )
iFixDataloss (Guo2022DetectingAF, ) Static Dynamic Fuzzing Search-based Testing Static Program Repair Detect Fix Activity  (Guo2022DetectingAFTool, )
RCHDroid (Chen2023TransparentRC, ) Dynamic Runtime Management Garbage Collection Fix Activity  (Chen2023TransparentRCTool, )
VALA (Lu2022DetectingRU, ) Static Dataflow Analysis Symbolic Execution Detect Activity Resource  (Lu2022DetectingRUTool, )
ACETON (Jabbarvand2020AutomatedCO, ) DL Dynamic Deep Learning Feature Engineering Runtime Monitoring Detect Resource  (Jabbarvand2020AutomatedCOTool, )
AsyncChecker(Pan2020StaticAC, ) Static Pattern Checking Interprocedural- Code Analysis Detect AsyncTask  (Pan2020StaticACTool, )
TAPIR (Li2019CharacterizingAD, ) Static Code Pattern Analysis Detect Image  (Li2019CharacterizingADTool, )
Lyu et al. (Lyu2018RemoveRF, ) Static Dataflow Analysis Program transformation Detect Fix Database Loop  (Lyu2018RemoveRFTool, )
ServDroid (Song2019ServDroidDS, ) Static Code Pattern Analysis Interprocedural- Code Analysis Detect Service  (Song2019ServDroidDSTool, )
Static Analysis. Static approaches are the art of reasoning about the behavior of programs without actually running them. It has been widely used (70.00%) to either detect or fix performance issues, as shown in Table 3. Specifically, taint analysis, symbolic execution, code pattern analysis and data flow analysis are the most commonly adopted techniques for rule-based static analysis as many performance issues arise from common code patterns. Additionally, static approaches, such as feature engineering, are used to assist deep learning approaches by extracting semantic code features (Kim2019AMD, ; Su2016ADL, ). Moreover, they can also provide fundamental information for dynamic fuzzing tools to trigger performance bugs on the fly (Guo2022DetectingAF, ). Furthermore, static approaches play an important role in fixing performance issues, particularly in generating patches for automatic program repair (Guo2022DetectingAF, ).
Dynamic Analysis. There are only four (i.e. 40.00%) primary studies that use dynamic analysis approaches to detect or fix performance bugs (especially for data loss issues). This result is align with expectation because dynamic analysis involves executing apps in a controlled environment and unable to cover all possible paths (dong2020timemachine, ). However, dynamic analysis is more resilient to evasion techniques (e.g., reflection (sun2021taming, ), obfuscation (sun2023demystifying, )) and can supplement static approaches to eliminate false positives (sun2022mining, ). Table 3 shows that three dynamic tools are designed to address data loss issues in the context of activity restarts associated with sophisticated lifecycle methods, which are challenging to be detected statically. We also discovered one study (Jabbarvand2020AutomatedCO, ) that employed dynamic monitoring to collect features for deep learning methodologies aimed at detecting energy issues caused by resource inefficiencies.
Deep Learning. With the increasing popularity of deep learning (DL) techniques, we are surprised to find that only one paper adopted DL techniques for performance analysis. Specifically, Jabbarvand et al. (Jabbarvand2020AutomatedCO, ) observed runtime features dynamically and leveraged DL models to locate energy anomalies. Although the number of DL-related studies is relatively small, DL techniques have been widely employed in various software engineering tasks, such as code analysis (Allamanis2017LearningTR, ), bug detection (Pradel2018DeepBugsAL, ), and program repair (zhao4254659autopatch, ). It is promising to see a garnered considerable research attention to employ ML/DL techniques to thwart Android performance issues.
Answers to RQ1.2: The majority of existing approaches primarily utilize static analysis (70.00%), with dynamic analysis (40.00%) coming next, and deep learning (10.00%). Only 10 out of the examined papers (20.83%) offer publicly accessible automated tools, highlighting a significant accessibility gap. Only 7 factors (33.33%) are addressed by existing automated tools, indicating incomplete coverage and potential oversight. Most tools prioritize issue detection over resolution, indicating a need for proactive problem-solving.
Table 4. A list of datasets used for Android mobile performance analysis.
Reference Year # APP Type Target Factor Link
Guo et al. (Guo2022DetectingAF, ) 2022 66 Real-World Activity  (Guo2022DetectingAFTool, )
Chen et al. (Chen2023TransparentRC, ) 2023 27 Real-World Activity  (Chen2023TransparentRCTool, )
Lu et al. (Lu2022DetectingRU, ) 2022 35 Real-World Activity Resource  (Lu2022DetectingRUTool, )
Jabbarvand et al. (Jabbarvand2020AutomatedCO, ) 2020 9 Micro Resource  (Jabbarvand2020AutomatedCOTool, )
AsyncBench (Pan2020StaticAC, ) 2020 69 Micro AsyncTask  (Pan2020StaticACTool, )
Li et al. (Li2019CharacterizingAD, ) 2019 19 Real-World Image  (Li2019CharacterizingADTool, )
Lyu et al. (Lyu2018RemoveRF, ) 2018 206 Real-World Database Operations Loop  (Lyu2018RemoveRFTool, )
Song et al. (Song2019ServDroidDS, ) 2019 45 Real-World Service  (Song2019ServDroidDSTool, )
Habchi et al. (Habchi2019TheRO, ) 2019 324 Real-World UI Rendering View Inefficient Method Hashmap  (Habchi2019TheROData, )
2.2.3. Result - Dataset
To evaluate the aforementioned automatic tools, researchers have proposed several datasets targeting unique performance factors. After examining all the public links or websites within the 48 identified studies, we were surprised to find only 9 (18.75%) unique open-sourced datasets, as shown in Table 4. Since most of these datasets are designed for diverse performance factors, it becomes infeasible to ensure a fair comparison. Additionally, the number of factors covered by existing datasets is quite limited, w.r.t 11 factors in total. The shortage of dataset suggests an urgent need of comprehensive dataset construction. Moreover, there are a limited number (i.e., 89 apps per dataset on average) of apps in each dataset, making it unrealistic to cover all potential performance issues. In other words, even though some tools might effectively detect issues, they would still suffer from soundness issues due to the existence of false negatives. Overall, more effort should be dedicated to creating more representative and comprehensive datasets.
Answers to RQ1.3: Available datasets covers only 11 out of 21 performance factors, highlighting a significant gap in tool evaluation. Existing datasets lack the representation needed to cover all performance issues adequately.
3. A Large-scale Exploration of Real-world Apps
In RQ1 (Section 2), we systematically summarised the Android performance issues in academia. However, there may still be gaps between cutting-edge studies and real-world scenarios. Thus, we conducted an explanatory study by examining real-world performance issues discussions. Specifically, we collected reported performance issues from three popular platforms of Android ecosystem: Google Play, Stack Overflow, and GitHub. We 1) reviewed user feedback on Google Play to assess the impact on end-users; 2) explored the discussions on Stack Overflow to gain insights into common issues, discussions, and solutions shared within the developer community; and 3) analysed issues and commit histories on GitHub to understand the challenges developers are facing. While detailing research gaps, we also present code patterns that we believe could inspire future works, contributing to our broader community.

### v2

@@ 1. Introduction
However, due to the non-functional nature of performance issues, their problems do not immediately manifest. Additionally, performance issues can be caused by multiple factors, requiring in-depth analysis and debugging to diagnose and identify the root causes (Liu2016FixingRL, ; Habchi2018OnAL, ; Jabbarvand2019SearchBasedET, ; Li2019CharacterizingAD, ; Liu2014GreenDroidAD, ). Furthermore, some performance issues only appear under specific usage scenarios or loads (Behrouz2016EnergyawareTM, ; Jabbarvand2019SearchBasedET, ; Yan2013SystematicTF, ; Farooq2018RuntimeDroidRR, ). Consequently, manually identifying performance issues is time-consuming and labor-intensive, making it difficult for developers to address all problems during the coding and testing phase. Therefore, it is essential for researchers to support developers in analyzing, detecting, and resolving Android performance issues (Habchi2018OnAL, ; Li2019DetectingAD, ; Li2022CombattingEI, ; Vsquez2015HowDD, ).
Our research fellows have proposed several approaches to mitigate the impact of performance issues (Li2019CharacterizingAD, ; Banerjee2014DetectingEB, ; Song2021IMGDroidDI, ; Rua2023ALE, ; Das2020CharacterizingTE, ; Vsquez2015HowDD, ; Hort2021ASO, ; Liu2019DroidLeaksAC, ; Habchi2019TheRO, ). For example, in the sense of static analysis, Li et al. (Li2019CharacterizingAD, ) developed TAPIR to analyze anti-patterns in app call graphs to detect inefficient image displays that can lead to memory consumption, energy consumption, and responsiveness issues. In addition, dynamic approaches, such as µDroid (Jabbarvand2017DroidAE, ), are proposed to detect energy waste due to resource under-utilization, wakelock misuse, and inefficient loops by monitoring abnormal energy consumption. However, given the scattered nature of current research in this field, it remains unclear whether performance issues have been comprehensively resolved. Specifically, it is unknown what the current state of Android app performance analysis is, whether current research is focused on key performance issues encountered in real-world settings, and to what extent existing methods effectively and comprehensively address these critical real-world challenges. Additionally, the capabilities of existing approaches are quite distinct, as confirmed by the fact of lacking a systematic taxonomy to map performance issues with their corresponding root causes (a.k.a, factors). Specifically, in this work, we formulate these concerns into two research questions (RQs) that we aim to answer through empirical evidence and explanatory results. The two RQs are summarized as follows.
RQ1: What are the primary performance issues that developers and users are most concerned with in the real-world settings (RQ1.1)? And what are the root causes (RQ1.2) and common code patterns (RQ1.3) of these issues?
RQ2: Do the performance issues investigated in academic research align with those that developers and users are most concerned with in real-world (RQ2.1)? To what extent do the tools developed in academic research address real-world performance issues (RQ2.2)? Are the datasets developed in academic research adequate for evaluating the effectiveness of these tools in tackling Android performance issues (RQ2.3)?
To answer these RQs, we began by conducting a large-scale analysis of real-world discussions on Android performance issues. Specifically, we collected 60,684 negative reviews from Google Play, 360,372 question threads from Stack Overflow, and 35,278 issues with 550,873 commits from GitHub to identify performance issues experienced by users and developers, analyze their root causes (i.e., contributing factors), and examine common code patterns. Following this, we conducted a systematic literature review on Android performance analysis to investigate whether existing works indeed touch on the leading performance issues encountered in real-world settings. Through a comparison of academic findings with real-world data, we developed a comprehensive taxonomy covering seven categories of performance issues and 72 contributing factors (Table 1). This taxonomy allowed us to identify substantial gaps in current academic research on Android performance analysis. Specifically, it turns out there is a substantial divergence in primary concerns among researchers, developers, and users. Additionally, it is surprising to find that among all identified factors, 57.14% have not been examined in academic research, while a substantial 76.39% remain unaddressed by existing tools, and 66.67% lack corresponding datasets. On top of these findings, we further propose future research directions to bridge these gaps.
In this paper, we make the following major contributions:
  • We conducted a large-scale exploratory study of real-world discussions on Android performance issues by analyzing crowd-sourced data from Google Play, Stack Overflow, and GitHub. This allowed us to identify the primary performance issues faced by users and developers, along with their root causes and common code patterns.
  • We conducted a systematic literature review on Android performance analysis to examine the performance issues that researchers focus on, the tools developed to address these issues, and the datasets used to evaluate the effectiveness of these tools.
  • By comparing literature with real-world data, we developed a taxonomy that categorizes seven types of performance issues and identifies 72 contributing factors, enabling us to assess how well current research objectives, tools, and datasets address real-world challenges. Based on these insights, we proposed future research directions to address the identified gaps.
2. A Large-scale Exploration of Real-world Discussions on Android Performance Issues
While performance issues frequently arise in real-world scenarios, it remains unclear which specific issues concern users and developers, along with their root causes and common code patterns. To address this, we conducted an exploratory study to collect, identify, and analyze real-world performance issues as discussed on popular online platforms.